OpenStudy (anonymous):

what is the multiplicative inverse of -1/9 ?

OpenStudy (shane_b):

Just flip the fraction...

OpenStudy (whpalmer4):

The multiplicative inverse of x is simply the number which when multiplied by x gives you 1.

OpenStudy (whpalmer4):

If x = -1/9, then -9/1 = -9 is the multiplicative inverse (another name is reciprocal). \[\frac{-1}{9}*\frac{-9}{1} = \frac{9}{9} = 1\]